OpenStudy (anonymous):

Please help?! Lisa can choose 4 of her 9 friends to come with her family on a camping trip. How many ways can she choose the friends that come with her? a. 126 b. 756 c. 3024 d. 6561 I think it is either a. 126 or c. 3024... is it something else? or what is the answer? I am confused. :/

Parth (parthkohli):

I know this one :D

Parth (parthkohli):

It'd be 9choose4 aka 9c4

OpenStudy (anonymous):

Can you please explain it to me too? I don't get it haha

Parth (parthkohli):

nCr = n!/[(n - r)! * r!]

OpenStudy (anonymous):

What does that mean? haha :P

Parth (parthkohli):

Lol

OpenStudy (anonymous):

yeah lol is right! haha I do not get these at all! lol (see what i did there?!) teepee

OpenStudy (anonymous):

***teeheeee lol not teepee.. auto correct is not on my side! haha

Parth (parthkohli):

9c4 = \(\Large \color{Black}{\Rightarrow {9! \over (9 - 4)! \times 4!} }\) \(\Large \color{Black}{\Rightarrow {9! \over 5! \times 4!} }\) \(\Large \color{Black}{\Rightarrow {9 \times 8 \times 7 \times 6 \over 4 \times 3 \times 2} }\)

Parth (parthkohli):

Lol teepee

Parth (parthkohli):

\(\Large \color{Black}{\Rightarrow {3024 \over 24} }\) \(\Large \color{Black}{\Rightarrow 126 }\)

OpenStudy (anonymous):

yeah hahaha :P Why do you stop at 6 though??

Parth (parthkohli):

I've cancelled 5!

OpenStudy (anonymous):

So the answer is a. 126 then? @ParthKohli ??

OpenStudy (anonymous):

oh and i see what you did there. I get it now haha.... @ParthKohli so the answer is 126 ways???

Parth (parthkohli):

Yes

Parth (parthkohli):

:P

Parth (parthkohli):

Do you have a scientific calculator?

OpenStudy (anonymous):

yes, but i do not know how to use it :P

Parth (parthkohli):

Use this site then, for example I've entered one already for you http://www.wolframalpha.com/input/?i=9c4

OpenStudy (anonymous):

oh wait no! i don't have one... haha nvm lol it's just a regular calculator! lol my bad

Parth (parthkohli):

:P

OpenStudy (anonymous):

Thanks for the link! It will be very useful! So is that a scientific calculator? ONe that is really easy to use right? teehee

Parth (parthkohli):

It's more than a scientific calculator! It has all subjects!

OpenStudy (anonymous):

OMG! it does! Wow! Once summer break starts, I'm gonna learn how to use it! or better yet, memorize how to use the entire thing!!! :) THanks for the link @ParthKohli and thanks for the help!

Parth (parthkohli):

No it automatically understands what you are trying to say!

OpenStudy (anonymous):

wow! THanks! ;) that's one handy dandy tool! how'd u find out about it?

Parth (parthkohli):

If you enter 9c4 or 9choose4 or 9 choose 4 or anything, it'll understand. It's very smart.

Parth (parthkohli):

It's famous all over the world. You didn't know?

OpenStudy (anonymous):

no i didn't haha really? it is? I had noooo idea lol

OpenStudy (anonymous):

thanks!! :) for the link AND the help!